OpenStudy (anonymous):

In a certain candy store, 3 pounds of candy and 2 pounds of mints cost $10.80, and 1 pound of candy and 3 pounds of mints cost $5.35. What is the cost per pound of the candy and the mints? If m = price of mints and c = price of candy, then which of the following expressions is equal to 5.35? 3m 1 + 3m c + 3m i got C

OpenStudy (anonymous):

C is correct

OpenStudy (texaschic101):

3c + 2m = 10.80 c + 3m = 5.35 -->(-3) ---------------- 3c + 2m = 10.80 -3c - 9m = - 16.05 (result of multiplying by -3) ----------------add -7m = - 5.25 m = -5.25/-7 m = 0.75 now sub 0.75 in for m 3c + 2m = 10.80 3c + 2(0.75) = 10.80 3c + 1.50 = 10.80 3c = 10.80 - 1.50 3c = 9.30 c = 9.30/3 c = 3.10 oops...I did a little too much.... answer is C. and by the way, mints cost 0.75 per lb and candy costs 3.10 per lb.